# Impact of positrons on electrical conductivity of hot and dense astrophysical plasma

### v1

@@ III.2 The electron-ion and positron-ion scattering matrix elements
where the electron, positron and ion 4-currents are given, respectively, by
@@
and $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of the currents transverse to $\bm{q}$. Here $\bm{p}_{24}=(\bm{p}_{2}+\bm{p}_{4})/2$, therefore, as $\bm{p}_{4}=\bm{p}_{2}+\bm{q}$, we obtain $\bm{p}_{24}^{\perp}=\bm{p}_{2}^{\perp}=\bm{p}_{4}^{\perp}\equiv\bm{p}^{\prime\perp}$. Squaring these matrix elements and performing the spin-averaging procedure yields the final result (see Ref. Harutyunyan2016 )
@@
$\hskip-28.45274pt+\frac{\bar{C}_{2}(\alpha,\theta)+2C^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{M^{2}_{i}|t_{\perp}|^{2}}\Biggr],$
@@
For the polarization tensor, we use the hard-thermal-loop expressions of QED, see Ref. Harutyunyan2016 Eqs. (40) and (41). In the regime of low-frequency scattering $(x=\omega/q\ll 1)$, relevant for electron-ion and positron-ion collisions, the longitudinal and transverse components (assuming ultra-relativistic electrons and positrons) reduce to
$\Pi_{L}(q,\omega)$ $=q_{D}^{2}\chi_{l},$ $\Pi_{T}(q,\omega)$ $=q_{D}^{2}\chi_{t},$ (72)
@@
$q_{D}^{2}=-\frac{4e^{2}}{\pi}\int_{0}^{\infty}d\epsilon\,p\,\epsilon\left(\frac{\partial f^{0+}}{\partial\epsilon}+\frac{\partial f^{0-}}{\partial\epsilon}\right).$ (73)
III.3 The collision rates
In this section, we summarize the results of computations of the collision rates given by Eqs. (39)–(42). The electron-ion collision rate was computed in Ref. Harutyunyan2016 with the assumption that the ion distribution is given in the classical Maxwell-Boltzmann form. This result can be generalized straightforwardly to obtain the electron-ion and positron-ion collision rates (39) and (40). We will use the simplified formula given by Eq. (32) of Ref. Harutyunyan2016 , which corresponds to the limit of static scattering. The final formulas read
@@ III.3 The collision rates
$\int_{0}^{2p}dq\,q^{3}S_{i}(q)F_{i}^{2}(q)\frac{4\epsilon^{2}-q^{2}}{|q^{2}+\Pi_{L}|^{2}}.$
@@
where the first and second terms represent the scattering and annihilation diagrams, respectively. Note that the variable $\epsilon$ stands for the energy of the initial state electron in both integrals, whereas the variable $\epsilon^{\prime}$ represents the energy of the final electron in the scattering integral and that of the initial positron in the annihilation integral. Thus, the variables $q$ and $\omega$ are the momentum and energy of the transferred plasmon in both cases, but the integration limits are different. The auxiliary functions in Eq. (75) are given by
@@
The lower bound for $\omega$ in the annihilation integral is $\omega_{\rm min}={\rm max}\{\epsilon,\epsilon^{\prime\prime}\}+m$, since the plasmon produced in the annihilation carries the total energy of the electron–positron pair.
Because the matrix element for $pe$ collisions is the same as the one for epep collisions, see Appendix A, we see from Eqs. (41) and (42) that the $pe$ collision rate can be obtained from the $ep$ collision rate by simply replacing $f^{0+}\leftrightarrow f^{0-}$ and $l^{-}\rightarrow l^{+}$, therefore
@@ IV Numerical results
We performed numerical calculations of the electrical conductivities from Eq. (49) in the regime relevant to neutron star outer crusts, considering a plasma composed of a single ion species, $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}, across the entire temperature–density range studied. Below, we first present the phase structure of crustal matter, followed by a discussion of our new conductivity results. For practical astrophysical applications, all numerical values are given in c.g.s. units.
@@ IV.1 Phase diagram of crustal plasma
Figure 3: The temperature-density phase diagram of crustal plasma composed of carbon. Ionic component of plasma forms a Boltzmann gas above the Coulomb temperature $T_{C}$, a classical liquid at $T_{p}\leq T\leq T_{C}$, and a quantum liquid below the plasma temperature $T_{p}$. Electrons become degenerate below the Fermi temperature $T_{F}$. The three curves around $T_{F}$ correspond to temperatures where the ratio of the total electron density $n^{-}$ to the net electron density $n_{e}$ reaches the value $r$.
To describe the state of the crustal plasma, we introduce the following physical quantities. The ion mass is defined as $M=Am_{n}$, where $A$ is the ion mass number, and $m_{n}\simeq 939$ MeV is the average nucleon mass. The ion number density $n_{i}$ is related to the density of matter $\rho$ by $n_{i}=\rho/M$, and we also define the radius of the spherical volume per ion (so-called the Wigner Seitz cell Shapiro1986 ) as $a=(4\pi n_{i}/3)^{-1/3}$.
@@
Electrons become degenerate below the Fermi temperature $T_{F}=\epsilon_{F}-m$, where the Fermi energy is defined as $\epsilon_{F}=(p_{F}^{2}+m^{2})^{1/2}$, and the Fermi momentum is given by $p_{F}=(3\pi^{2}n_{e})^{1/3}$. At finite temperatures, once the thermal energy of matter becomes sufficient to allow for creation of electron-positron pairs, a thermal population of positrons appears. The electron number excess over positrons $n_{e}$ is now the relevant quantity that determine the charge neutrality condition $n_{e}=Zn_{i}$. The electron chemical potential is found from the value implied by the charge conservation for electron excess $n_{e}=n^{-}-n^{+}$ for any given values of $\rho$ and $T$. The positron/electron proper number densities are given by
@@
where the $f^{0\pm}$ are the Fermi distributions of electron and positrons introduced earlier. Clearly, $n^{-}>n_{e}$ when positrons’ population is not negligible.
@@
In order to assess the relative importance of positrons in the plasma, we define the ratio $r=n^{-}/n_{e}$. Physically, $r$ is expected to rise rapidly with the temperature as a result of the fast opening of kinematic phase space for the pair creation with increasing $T$. This feature is clearly seen in Fig. 3, where we have plotted three lines which correspond to the values $r=2,5,10$. These curves highlight the crucial importance of including electron–positron pair creation processes in the composition of crustal matter, even at temperatures near the Fermi temperature. The ratio $r$ also increases with decreasing density at fixed $T$, as the pair creation rate is closely related to the ratio $T/T_{F}$. The shaded region at the top of the phase diagram corresponds to the regime where the number of created positrons, $n^{+}$, already exceeds the number of electrons in the cold $(T\ll T_{F})$ plasma. At temperatures $T\gg T_{F}$, the number of pairs increases approximately cubically with temperature, as implied by Eq. (89), yielding $n^{+}\simeq n^{-}\gg n_{e}$ in this regime.
IV.2 Electrical conductivity from electron-ion and positron-ion collisions
@@
For clarity, we first consider the conductivities that include only electron–ion and positron–ion collisions. The full conductivities, which also account for electron–positron collisions, will be addressed in the next subsection. As indicated by Eqs. (43) and (74), the relaxation times for electrons and positrons are essentially identical in the regime where positrons are sufficiently abundant, since the integrals over combintations of Fermi functions in the numerators and denominators largely cancel. Consequently, the primary distinction between the electron and positron contributions to the electrical conductivity arises from the Fermi distribution functions in Eq. (49).
@@
In the opposite limit of high temperatures $T\gg T_{F}$, we have also $T\gg\mu$, electrons and positrons are ultrarelativistic with the average thermal energy $\bar{\epsilon}\simeq 3T$, and their number densities are almost equal $n^{+}\simeq n^{-}\equiv\bar{n}\sim\bar{\epsilon}^{3}$. In this case $f^{0+}\simeq f^{0-}\equiv f^{0}\ll 1$, and Eq. (49) can be approximated as
@@
Thus, the formulas for conductivity in both – strongly degenerate and nondegenerate regimes have the same form, but contain different energy scales, specifically $\epsilon_{F}$ in the degenerate regime and $\bar{\epsilon}\simeq 3T$ in the nondegenerate, ultrarelativistic regime.
@@
The relaxation time (74) in the physical regimes of interest was discussed in detail in Ref. Harutyunyan2016 . On average, it was found that for carbon, $\tau$ varies with density and temperature according to the scaling $\tau(\epsilon)\propto\epsilon^{2}\rho^{-0.9}T^{-0.2}$. Using this scaling and the rough estimates $n_{e}\sim\epsilon_{F}^{3}$ and $\bar{n}\sim\bar{\epsilon}^{3}$, Eqs. (90) and (91) yield a universal scaling $\sigma\propto\epsilon^{4}\rho^{-0.9}T^{-0.2}$. This corresponds approximately to $\sigma\propto\rho^{0.5}T^{-0.2}$ in the degenerate regime and $\sigma\propto\rho^{-1}T^{4}$ in the nondegenerate regime.
Figure 4: Dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the temperature indicated in the plot. These are the conductivities resulting only from eiei and pipi collisions. The dotted lines show the conductivity from Ref. Harutyunyan2016 where positrons’ abundance in the matter is neglected.
Figure 5: The temperature dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the density indicated in the plot. These are the conductivities resulting only from eiei and pipi collisions. The dotted lines show the conductivity from Ref. Harutyunyan2016 where positrons’ abundance in the matter is neglected.
@@
Figure 4 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$, as well as the total conductivity $\sigma=\sigma^{-}+\sigma^{+}$ as functions of density for three temperature values, which are chosen to cover the range from the degenerate regime ($T=0.1$ MeV) to the nondegenerate regime ($T=10$ MeV). The intermediate value $T=1$ MeV is a representative of the transition between these regimes, which occurs around $\rho\simeq 10^{8}$ g cm-3, see Fig. 3. For comparison, we also plot the conductivities obtained previously in Ref. Harutyunyan2016 , where the presence of positrons in the matter was not taken into account.
@@
The qualitative behavior of the conductivity in the degenerate regime is opposite to that in the transitional and nondegenerate regimes. This contrast can be understood by noting that at a temperature of $T=0.1$ MeV, positrons are effectively absent, since the temperature is far below the electron rest mass. The positron distribution is also strongly suppressed at $\rho\geq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV due to Pauli blocking: as the Fermi energy increases, the number of kinematically accessible states for pair creation diminishes. In both situations, the present results for $\sigma$ coincide with the earlier work Harutyunyan2016 , as expected. In this regime, the conductivity scales as $\sigma\propto\rho^{0.5}$, because the increase in the electron density of states with $\rho$ outweighs the decrease in $\tau_{F}$.
@@
Positrons are seen to make a significant contribution to the conductivity in the low-density regime $\rho\leq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV, and across nearly the entire outer crust ($\rho\leq 10^{11}\,\mathrm{g\,cm^{-3}}$) for $T=10$ MeV. In this regime, the behavior of the conductivity reverses: at fixed temperature, it decreases with density, following the power law $\sigma\propto\tau\propto\rho^{-1}$. This occurs because the total number of electron–positron pairs, $\bar{n}$, is determined by temperature alone and is independent of the matter density.
@@
The two conductivity branches representing the degenerate and nondegenerate regimes merge smoothly at a minimum located at their intersection. This transition point corresponds approximately to the temperature $T\simeq T_{F}/3$, where the two characteristic energy scales become comparable.
@@
Let us now turn to the temperature dependence of the conductivity (see Fig. 5). The low-temperature branches again reproduce the results of Ref. Harutyunyan2016 , decreasing slowly with temperature due to the similar behavior of $\tau_{F}$, which is governed by ion–ion correlations. In Ref. Harutyunyan2016 , the conductivity was found to reach a minimum at the transition point $T\simeq T_{F}/3$, resulting from the reversed temperature dependence of the relaxation time in the nondegenerate regime.

### v2

@@ III.2 Electron-ion and positron-ion scattering matrix elements
where the electron, positron, and ion 4-currents are given, respectively, by
@@
and $\bm{J}_{\perp},\bm{J}^{\prime}_{\perp}$ are the components of the currents transverse to $\bm{q}$. Here, $\bm{p}_{24}=(\bm{p}_{2}+\bm{p}_{4})/2$; therefore, as $\bm{p}_{4}=\bm{p}_{2}+\bm{q}$, we obtain $\bm{p}_{24}^{\perp}=\bm{p}_{2}^{\perp}=\bm{p}_{4}^{\perp}\equiv\bm{p}^{\prime\perp}$. Squaring these matrix elements and performing the spin-averaging procedure yields the final result (see Ref. Harutyunyan2016 ),
@@
$\hskip-28.45274pt+\frac{\bar{C}_{2}(\alpha,\theta)+2C^{2}_{1}(\alpha,\theta)\cos^{2}\varphi}{M^{2}_{i}|t_{\perp}|^{2}}\Biggr],\quad$
@@
For the polarization tensor, we use the hard-thermal-loop expressions of QED; see Ref. Harutyunyan2016 , Eqs. (43) and (44). This approximation assumes that $\omega\ll\epsilon$ and $q\ll p$; i.e., the 4-momentum of the plasmon is much smaller than that of the electron or positron inside the fermionic loop. Numerical tests show that in the whole range of densities and temperatures studied here, we effectively have $\omega\lesssim q\ll p\simeq\epsilon$; therefore, the hard-thermal-loop approximation to the polarization tensor works very well. In the regime of low-frequency scattering $(x=\omega/q\ll 1)$, relevant for electron-ion and positron-ion collisions, the longitudinal and transverse components (assuming ultrarelativistic electrons and positrons) reduce to
$\Pi_{L}(q,\omega)=q_{D}^{2}\chi_{l},\quad\Pi_{T}(q,\omega)=q_{D}^{2}\chi_{t},$ (72)
@@
$q_{D}^{2}=-\frac{4e^{2}}{\pi}\int_{m}^{\infty}d\epsilon\,p\,\epsilon\left(\frac{\partial f^{0+}}{\partial\epsilon}+\frac{\partial f^{0-}}{\partial\epsilon}\right).$ (73)
III.3 Collision rates
In this section, we summarize the results of the computations of the collision rates given by Eqs. (39)–(42). The electron-ion collision rate was computed in Ref. Harutyunyan2016 with the assumption that the ion distribution is given in the classical Maxwell-Boltzmann form. This result can be generalized straightforwardly to obtain the electron-ion and positron-ion collision rates (39) and (40). We will use the simplified formula given by Eq. (32) of Ref. Harutyunyan2016 , which corresponds to the limit of static scattering (this approximation is well justified for ions with $Z_{i}>1$). The final formulas read
@@ III.3 Collision rates
$\times\int_{0}^{2p}dq\,q^{3}S_{i}(q)F_{i}^{2}(q)\frac{4\epsilon^{2}-q^{2}}{|q^{2}+\Pi_{L}|^{2}}.\quad$
@@
where the first and second terms represent the scattering and annihilation diagrams, respectively. Note that the variable $\epsilon$ stands for the energy of the initial state electron in both integrals, whereas the variable $\epsilon^{\prime}$ represents the energy of the initial positron in the scattering integral and that of the final electron in the annihilation integral. Thus, the variables $q$ and $\omega$ are the momentum and energy of the transferred plasmon in both cases, but the integration limits are different. The auxiliary functions in Eq. (75) are given by
@@
The lower bound for $\omega$ in the annihilation integral is $\omega_{\rm min}={\rm max}\{\epsilon,\epsilon^{\prime}\}+m$ since the plasmon produced in the annihilation carries the total energy of the electron–positron pair.
Because the matrix element for $pe$ collisions is the same as the one for $ep$ collisions (see Appendix A), we see from Eqs. (41) and (42) that the $pe$ collision rate can be obtained from the $ep$ collision rate by simply replacing $f^{0+}\leftrightarrow f^{0-}$ and $l^{-}\rightarrow l^{+}$; therefore,
@@ IV Numerical results
We performed numerical calculations of the relaxation times using Eqs. (33), (34) and (39)–(42), and the electrical conductivities from Eq. (49) in the regime relevant to neutron star outer crusts, considering a plasma composed of a single ion species across the entire temperature–density range studied. Below, we first present the phase structure of crustal matter, followed by a discussion of our new results for relaxation times and conductivities. A more detailed discussion including the phase structure of the crust and relaxation times is provided for $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$} nuclei, but the final results for conductivities are also presented for $^{\makebox[7.97226pt][r]{$\scriptstyle 56$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Fe}$} and $^{\makebox[3.98613pt][r]{$\scriptstyle 4$}}_{\makebox[3.98613pt][r]{$\scriptstyle$}}\mathrm{He}$} nuclei. For practical astrophysical applications, all numerical values are given in in centimeter-gram-second units.
@@ IV.1 Phase diagram of crustal plasma
Figure 3: Temperature-density phase diagram of crustal plasma composed of carbon. The ionic component of plasma forms a Boltzmann gas above the Coulomb temperature $T_{C}$, a classical liquid at $T_{p}\leq T\leq T_{C}$, and a quantum liquid below the plasma temperature $T_{p}$. Electrons become degenerate below the Fermi temperature $T_{F}$. The three curves around $T_{F}$ correspond to temperatures where the ratio of the total electron density $n^{-}$ to the net electron density $n_{e}$ reaches the value $r$.
Figure 4: Electron chemical potential $\mu_{e}$ (panel a) and the Debye wave number $q_{D}$ (panel b) as functions of the temperature for three values of the density indicated in the plot. The dotted lines show the same quantities in the case where positrons are neglected.
To describe the state of the crustal plasma, we introduce the following physical quantities. The ion mass is defined as $M=Am_{n}$, where $A$ is the ion mass number and $m_{n}\simeq 939$ MeV is the average nucleon mass. The ion number density $n_{i}$ is related to the density of matter $\rho$ by $n_{i}=\rho/M$, and we also define the radius of the spherical volume per ion (the so-called Wigner Seitz cell Shapiro1983 ) as $a=(4\pi n_{i}/3)^{-1/3}$.
@@
Figure 5: Dependence of electron (solid lines) and positron (dashed lines) relaxation times on density for the temperature fixed at (a) $T=1$ MeV; and (b) $T=10$ MeV. The dotted lines show the electron-ion relaxation time $\tau_{ei}=\epsilon/\nu_{ei}$, the dash-dotted lines show the electron-positron relaxation time $\tau_{ep}=\epsilon/2\nu_{ep}$, and the dash–double-dotted lines show the positron-electron relaxation time $\tau_{pe}=\epsilon/2\nu_{pe}$.
Electrons become degenerate below the Fermi temperature $T_{F}=\epsilon_{F}-m$, where the Fermi energy is defined as $\epsilon_{F}=(p_{F}^{2}+m^{2})^{1/2}$, and the Fermi momentum is given by $p_{F}=(3\pi^{2}n_{e})^{1/3}$. At finite temperatures, once the thermal energy of matter becomes sufficient to allow for the creation of electron-positron pairs, a thermal population of positrons appears. The electron number excess over positrons, $n_{e}$ is now the relevant quantity that determines the charge neutrality condition $n_{e}=Zn_{i}$. The electron chemical potential is found from the value implied by the charge conservation for electron excess $n_{e}=n^{-}-n^{+}$ for any given values of $\rho$ and $T$. The positron and electron proper number densities are given by
@@
where $f^{0\pm}$ are the Fermi distributions of electrons ()(-) and positrons $(+)$ introduced earlier. Clearly, $n^{-}>n_{e}$ when the population of positrons is not negligible.
@@
In order to assess the relative importance of positrons in the plasma, we define the ratio $r=n^{-}/n_{e}$. Physically, $r$ is expected to rise rapidly with the temperature as a result of the fast opening of kinematic phase space for the pair creation with increasing $T$. This feature is clearly seen in Fig. 3, where we have plotted three lines which correspond to the values $r=2,5,10$. These curves highlight the crucial importance of including electron–positron pair creation processes in the composition of crustal matter, even at temperatures near the Fermi temperature. The ratio $r$ also increases with decreasing density at fixed $T$, as the pair creation rate is closely related to the ratio $T/T_{F}$. The shaded region at the top of the phase diagram corresponds to the regime where the number of created positrons (or, equivalently, that of created pairs), $n^{+}$, already exceeds the number of electrons in the cold $(T\ll T_{F})$ plasma. At temperatures $T\gg T_{F}$, the number of pairs increases approximately cubically with temperature, as implied by Eq. (89), yielding $n^{+}\simeq n^{-}\gg n_{e}$ in this regime.
We also show the electron chemical potential $\mu_{e}$ and the Debye wave number in Fig. 4 as functions of temperature for three values of the density. The dotted lines correspond to the case in which the positron abundance is neglected. In the degenerate limit, $T\ll T_{F}$, the positron contribution vanishes, and we recover $\mu_{e}\simeq\epsilon_{F}$ and $q_{D}^{2}\simeq 4e^{2}p_{F}\epsilon_{F}/\pi$, which depend only on the density, as seen in the figure. In the opposite, high-temperature limit $T\gg T_{F}$, the chemical potential tends to zero, $\mu_{e}\to 0$, and electrons and positrons contribute nearly equally to the Debye wave number. In this regime, Eq. (73) yields the approximation $q_{D}^{2}\simeq 16e^{2}T^{2}/\pi.$ This behavior contrasts with the high-temperature scaling $q_{D}^{2}\simeq 4e^{2}\pi n_{e}/T$ obtained when the positron contribution is neglected. Consequently, the presence of positrons leads to significantly larger, density-independent Debye wave numbers and therefore to stronger screening of scattering amplitudes compared to the positron-free case.
IV.2 Relaxation times
Figure 6: Temperature dependence of electron (solid lines) and positron (dashed lines) relaxation times for the density fixed at (a) $\rho=10^{6}$ g cm-3 and (b) $\rho=10^{8}$ g cm-3. The dotted lines show the electron-ion relaxation time $\tau_{ei}=\epsilon/\nu_{ei}$, the dash-dotted lines show the electron-positron relaxation time $\tau_{ep}=\epsilon/2\nu_{ep}$, and the dash–double-dotted lines show the positron-electron relaxation time $\tau_{pe}=\epsilon/2\nu_{pe}$.
In this section, we discuss the electron and positron relaxation times arising from electron-ion, positron-ion, and electron-positron scattering processes. Let us recall that the electron–positron and positron–electron collision rates given by Eqs. (75) and (86) contain two distinct microscopic interaction channels: ordinary scattering and pair annihilation–creation processes. The latter is found to contribute significantly less to the total collision rate than the genuine scattering channel.
Figure 5 shows the density dependence of the relaxation times at fixed temperatures $T=1$ MeV (left panel) and $T=10$ MeV (right panel). Since the relaxation times are energy dependent $(\tau^{\pm}=\epsilon X^{\pm}$ with $X^{\pm}=\mathrm{const})$, their evaluation depends on the degeneracy of the electrons. To distinguish between the degenerate and nondegenerate regimes, we introduce the characteristic temperature $T^{*}\equiv T_{F}/3$, which serves as the transition scale between the two regimes. In degenerate matter, $T\leq T^{*}$, the electron relaxation time $\tau^{-}$ is evaluated at the electron Fermi energy $\epsilon_{F}$, while in the nondegenerate ultrarelativistic regime, $T\geq T^{*}$, it is evaluated at the thermal energy $\bar{\epsilon}=3T$. The positron relaxation time $\tau^{+}$, by contrast, is always evaluated at the thermal energy.
In order to assess the relative importance of the separate scattering processes, we show the partial contributions to the relaxation times due to electron-ion and electron-positron collisions. At high densities, where the positrons are irrelevant, the electron relaxation times coincide with those due to electron-ion collisions $\tau_{ei}\equiv\epsilon/\nu_{ei}$. In this regime $\tau^{-}$ decreases with density because of increasing electron-ion scattering rates; here, we recover the old results of Ref. Harutyunyan2016 .
In the opposite limit of dilute matter, where $T\gg T_{F}$, electron-positron collisions strongly dominate $ei$ and $pi$ collisions, and $\tau^{-}$ and $\tau^{+}$ are close to their limiting values $\tau_{ep}\simeq\tau_{pe}\equiv\epsilon/2\nu_{ep}$, as follows from Eq. (38). Here $\tau_{ep}$ ($\tau_{pe}$) tends to a density-independent value below the transition point, but increases (decreases) above that point, as the positron population drops; the net electron density increases when matter enters the degenerate regime.
The total electron relaxation time $\tau^{-}$ interpolates between these two regimes, approaching a density-independent value $\tau_{ep}$ at small densities and decreasing at higher densities. The transition point moves to higher densities for larger values of $T$. We remark also that the longitudinal and transverse parts of the scattering amplitudes contribute almost equally to the electron-positron collision rates $\nu_{ep}$ and $\nu_{pe}$ in the regime where these are relevant.
The behavior of the positron relaxation time $\tau^{+}$ is very similar to that of $\tau^{-}$ but with slightly lower values. The partial contribution $\tau_{pi}$ is very close to $\tau_{ei}$ in the regime where it is relevant; therefore, it is not shown in the figures. Indeed, as indicated by Eqs. (43) and (74), the electron-ion and positron-ion collision rates are essentially identical in the regime where positrons are sufficiently abundant, since the integrals over combinations of Fermi functions in the numerators and denominators largely cancel.
Figure 6 shows the temperature dependence of the relaxation times for several fixed values of the density. We see again that, in the strongly degenerate regime $T\ll T_{F}$, we have $\tau^{-}\simeq\tau_{ei}$, which decreases with the temperature due to the ion structure factor $S(q)$ Harutyunyan2016 . At intermediate temperatures $T\sim T^{*}$, where the matter is semidegenerate, $\tau_{ei}$ changes its behavior from a decreasing to an increasing function. This transition reflects the approximate scaling $\tau_{ei}\propto\epsilon\propto T$. In the semidegenerate regime, there exists a temperature window in which electron–positron collisions remain subdominant, i.e., $\tau_{ei}\gtrsim\tau_{ep}$. As a result, the effective electron relaxation time $\tau^{-}$ increases in this domain. However, at higher temperatures $T\gg T_{F}$, the electron-positron collisions become dominant, and we have $\tau^{-}\simeq\tau_{ep}\propto T^{-1}$, which follows from dimensional arguments. Thus, as a result of nontrivial convolution of electron-ion and electron-positron collisions, the electron relaxation time $\tau^{-}$ shows a nonmonotonic behavior with temperature, being a decreasing function in both limiting cases (with a higher slope in the nondegenerate regime), and increasing in the transitional regime which interpolates between these two limiting cases. The behavior of the positron relaxation time $\tau^{+}$ is very similar to that of $\tau^{-}$ in the transitional and nondegenerate regimes where it is relevant, albeit with slightly smaller values.
IV.3 Electrical conductivity from electron-ion and positron-ion collisions
Figure 7: Dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the temperature indicated in the plot. These are the conductivities resulting only from eiei and pipi collisions. The dotted lines show the conductivity from Ref. Harutyunyan2016 where positron contribution is neglected.
Figure 8: Temperature dependence of electronic (dashed lines) and positronic (dash-double-dotted lines) partial conductivities and their sum (solid lines) for three values of the density indicated in the plot. These are the conductivities resulting only from eiei and pipi collisions. The dotted lines show the conductivity from Ref. Harutyunyan2016 where positron contribution is neglected.
@@ IV.3 Electrical conductivity from electron-ion and positron-ion collisions
For clarity, we first consider the conductivities that include only electron–ion and positron–ion collisions. The full conductivities, which also account for electron–positron collisions, will be addressed in the next subsection. As discussed above, $\tau_{ei}\simeq\tau_{pi}$ in the regime where positrons are sufficiently abundant; therefore, the primary distinction between the electron and positron contributions to the electrical conductivity arises from the Fermi distribution functions in Eq. (49).
@@
In the opposite limit of high temperatures $T\gg T_{F}$, we also have $T\gg\mu_{e}$, electrons and positrons are ultrarelativistic with the average thermal energy $\bar{\epsilon}\simeq 3T$, and their number densities are almost equal, $n^{+}\simeq n^{-}\equiv\bar{n}\sim\bar{\epsilon}^{3}$. In this case, $f^{0+}\simeq f^{0-}\equiv f^{0}\ll 1$, and Eq. (49) can be approximated as
@@
Thus, the formulas for conductivity in both strongly degenerate and nondegenerate regimes have the same form but contain different energy scales, specifically, $\epsilon_{F}$ in the degenerate regime and $\bar{\epsilon}\simeq 3T$ in the nondegenerate, ultrarelativistic regime.
@@
The relaxation time (74) in the physical regimes of interest was discussed in detail in Ref. Harutyunyan2016 . On average, it was found that, for carbon, $\tau$ varies with density and temperature according to the scaling $\tau(\epsilon)\propto\epsilon^{2}\rho^{-0.9}T^{-0.2}$. Using this scaling and the rough estimates $n_{e}\sim\epsilon_{F}^{3}$ and $\bar{n}\sim\bar{\epsilon}^{3}$, Eqs. (90) and (91) yield a universal scaling $\sigma\propto\epsilon^{4}\rho^{-0.9}T^{-0.2}$. This corresponds approximately to $\sigma\propto\rho^{0.5}T^{-0.2}$ in the degenerate regime and $\sigma\propto\rho^{-1}T^{4}$ in the nondegenerate regime.
@@
Figure 7 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$, as well as the total conductivity $\sigma=\sigma^{-}+\sigma^{+}$ as functions of density for three temperature values, which are chosen to cover the range from the degenerate regime ($T=0.1$ MeV) to the nondegenerate regime ($T=10$ MeV). The intermediate value $T=1$ MeV is a representative of the transition between these regimes, which occurs around $\rho\simeq 10^{8}$ g cm-3; see Fig. 3. For comparison, we also plot the conductivities obtained previously in Ref. Harutyunyan2016 , where the presence of positrons in matter was not taken into account.
Figure 9: Same as Fig. 7, but with the inclusion of epep collisions.
Figure 10: Same as Fig. 8, but with the inclusion of epep collisions.
@@
The qualitative behavior of the conductivity in the degenerate regime is opposite to that in the transitional and nondegenerate regimes. This contrast can be understood by noting that at a temperature of $T=0.1$ MeV, positrons are effectively absent since the temperature is far below the electron rest mass. The positron distribution is also strongly suppressed at $\rho\geq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV due to Pauli blocking: As the Fermi energy increases, the number of kinematically accessible states for pair creation diminishes. In both situations, the present results for $\sigma$ coincide with the earlier work Harutyunyan2016 , as expected. In this regime, the conductivity scales as $\sigma\propto\rho^{0.5}$, because the increase in the electron density of states with $\rho$ outweighs the decrease in $\tau_{F}$.
@@
Positrons are seen to make a significant contribution to the conductivity in the low-density regime $\rho\leq 10^{8}\,\mathrm{g\,cm^{-3}}$ for $T=1$ MeV, and across nearly the entire outer crust ($\rho\leq 10^{11}\,\mathrm{g\,cm^{-3}}$) for $T=10$ MeV. In this regime, the behavior of the conductivity reverses: At fixed temperature, it decreases with density, following the power law $\sigma\propto\tau\propto\rho^{-1}$. This occurs because in this regime, the total number of electron–positron pairs, $\bar{n}$, is determined by temperature alone and is independent of the matter density.
@@
The two conductivity branches representing the degenerate and nondegenerate regimes merge smoothly at a minimum located at their intersection. This transition point corresponds approximately to the temperature $T\simeq T^{*}$, where the two characteristic energy scales become comparable.
@@
Let us now turn to the temperature dependence of the conductivity (see Fig. 8). The low-temperature branches again reproduce the results of Ref. Harutyunyan2016 , decreasing slowly with temperature due to the similar behavior of $\tau_{F}$, which is governed by ion–ion correlations. In Ref. Harutyunyan2016 , the conductivity was found to reach a minimum at the transition point $T\simeq T^{*}$, resulting from the reversed temperature dependence of the relaxation time in the nondegenerate regime.